OpenStudy (anonymous):

Solve 2|5x − 2| + 4 = 18 for x. Show your work to receive full credit.

OpenStudy (anonymous):

2|5x-2|+4=18 -4 -4 2|5x-2|=14 *2 *2 |5x-2|=28 5x-2=28 5x-2=-28 +2 +2 +2 +2 5x=30 5x=-25 5x=6 5x=-5